Functions f , g : R → R are defined, respectively, by f (x) = x2 + 3x + 1, g (x) = 2x – 3,
find  f o g  

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Given that,
f(x)=x^2+3x+1
g(x)=2x–3

fog = f(g(x))

= f(2x - 3)
= (2x-3)^2+3(2x-3) + 1
= 4x^2+9–12x + 6x–9 + 1
= 4x^2–6x+1

Thanks